### v2

# Adapting Large Language Models for Education: Foundational Capabilities, Potentials, and Challenges

Abstract
Online education platforms, leveraging the internet to distribute education resources, seek to provide convenient education but often fall short in real-time communication with students. They often struggle to offer personalized education resources due to the challenge of addressing the diverse obstacles students encounter throughout their learning journey. Recently, the emergence of large language models (LLMs), such as ChatGPT, offers the possibility for resolving this issue by comprehending individual requests. Although LLMs have been successful in various fields, creating an LLM-based education system is still challenging for the wide range of educational skills required. This paper reviews the recently emerged LLM researches related to educational capabilities, including mathematics, writing, programming, reasoning, and knowledge-based question answering, with the aim to explore their potential in constructing the next-generation intelligent education system. Based on the current development status, we further outline two approaches for an LLM-based education system: a unified approach and a mixture-of-expert (MoE) approach. Finally, we explore the challenges and future directions, providing new research opportunities and perspectives on adapting LLMs for education.
1 Introduction
Education plays a vital role in shaping individuals’ futures as it forms the foundation for providing people with knowledge, skills, and critical thinking abilities. Conventional education systems heavily rely on teachers for imparting knowledge to students, which place a significant demand on educational resources. However, the advent of online education has substantially lowered the cost of accessing these educational materials. Many people are conveniently acquiring knowledge through online courses and exercises. However, online education faces the issue of not being able to interact with students in real-time like a teacher. Each student inevitably encounters unique challenges during their learning journey. Presently, online education platforms typically offer static course videos and exercises, leaving students unable to ask questions or seek dynamic solutions to their specific issues. Therefore, developing a teaching assistant model capable of addressing students’ individual concerns represents a crucial step forward in advancing online education.
Figure 1: An example of LLM-based education systems integrating multiple abilities to solve student problems.
The emergence of large language models (LLMs) instills optimism in creating an intelligent education system. Since the launch of ChatGPT, LLMs have shown exceptional ability in understanding human knowledge and have been widely applied in various professional fields, including recommendation systems (Lin et al., 2023), healthcare (Liu et al., 2023d), economics (Li et al., 2023), and others. However, the challenge intensifies when constructing an LLM-based education system due to the broad spectrum of knowledge encompassed by education, including mathematics, reasoning, writing, and beyond.
The act of LLMs answering students’ questions can be considered as a process that involves using multiple educational capabilities. For instance, as depicted in Figure 1, the question posed by the student requires the education assistant to possess both mathematical and programming skills simultaneously. Therefore, summarizing the development status of LLMs’ education-related capabilities is meaningful for building the next-generation intelligent education system. In this paper, we study LLMs from the perspective of these education-related capabilities and explore the potential of an LLM-based education system, aiming to provide valuable insights into the areas that need further improvement. We present our primary investigation in Figure 2. Our specific focus centers on five distinct capabilities of LLMs:
  • Mathematics: We primarily summarize the current development of LLMs in solving mathematical problems of varying difficulty and types, encompassing basic arithmetic operations, challenging problems, multi-modal problems, and math proof.
  • Writing: We investigate the performance of LLMs on a number of representative writing tasks to outline the problems LLMs face and potential future directions.
  • Programming: In accordance with human programming conventions, we divide LLMs’ programming process into two stages: code writing and code refinement. We review the researches in this area and summarize the remaining problems for code education.
  • Reasoning: We explore the capability of LLMs to perform reasoning in various ways, including supervised fine-tuning, prompt engineering, and hybrid methods, and explore their potential applications in the field of education.
  • Knowledge-based Question Answering: We investigate the development of LLMs in open-domain and domain-specific knowledge-based question answering. We hope to offer a comprehensive view of incorporating such capabilities into the education system.
Besides summarizing the current development status of various foundational abilities related to education, we propose two potential approaches for forming an LLM-based education system. One approach involves training a comprehensive language model with multiple capabilities, while the other is based on an LLM controller in a mixture-of-experts framework. We analyze the implementation difficulties and pros and cons of these two approaches.
LLMs hold vast potential in the field of education, and our work aims to offer insights for building an LLM-based intelligent education system. While Gan et al. (2023) anticipated the prospects of LLMs for education by exploring the various roles they can play in the education process, our work differs. We review the evolution of LLMs from the perspective of education-related capabilities. Additionally, we provide foresight into constructing feasible frameworks for LLM-based education systems. Our work emphasizes a comprehensive understanding of LLMs’ educational competencies and explores frameworks that can effectively integrate these abilities into the educational landscape.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.6em,font=, where level=2text width=5.3em,font=, where level=3text width=9.7em,font=, [Foundational
Capabilities,text width=6.3em [Mathematics, text width=4em [Basic Arithmetic Problems, text width=8em [Goat (Liu and Low, 2023), GENBERT (Geva et al., 2020),leaf, text width=17.5em] ] [Challenging Mathematics Problems,text width=10.3em [WizardMath (Luo et al., 2023a), MeatMath (Yu et al., 2023),leaf, text width=18.5em] ] [Multimodal Problems,text width=6.3em [Mathvista (Lu et al., 2023), Unigeo (Chen et al., 2022b),leaf, text width=17.3em] ] [Mathematical Proof,text width=6em [LeanDojo (Yang et al., 2023b), LEGO-Prover (Xin et al., 2023),leaf, text width=19.5em] ] ] [Writing,text width=2.5em [Text Summarization,text width=6.3em [BRIO (Liu et al., 2022), PROM (Ma et al., 2023b),leaf, text width=15.4em] ] [Grammatical Error Correction,text width=9.5em [CoEdit (Liu et al., 2022), GrammarGPT (Fan et al., 2023),leaf, text width=18em] ] ] [Programming,text width=4.1em [Code Writing, text width=4.8em [CodeLlama (Rozière et al., 2023), WizardCoder (Luo et al., 2023b),leaf, text width=20em] ] [Code Refinement, text width=6em [Codet (Chen et al., 2022a), SEIDR (Liventsev et al., 2023),leaf, text width=18.5em] ] ] [Reasoning, text width= 3em [Supervised Finetuning,text width=6.3em [CAGE (Rajani et al., 2019), Scratchpad (Nye et al., 2021),leaf, text width=18em] ] [Prompt Engineering,text width=6.3em [CoT (Wei et al., 2022), SelfConsistency (Wang et al., 2022),leaf, text width=18.5em] ] [Hybrid Methodologies,text width=6.3em [FineTuneCoT (Ho et al., 2022), STaR (Zelikman et al., 2022),leaf] ] ] [Knowledge-based QA , text width= 6.3em [Open-domain QA [Retro (Borgeaud et al., 2022), FRESHPROMPT (Vu et al., 2023),leaf, text width=19.8em] ] [Domain-specific QA,text width=6.5em [Reta-llm (Liu et al., 2023c), LeanContext (Arefeen et al., 2023),leaf]] ] ] ] ]
Figure 2: Summary of LLM’s education-related foundational capabilities.
2 Foundational Capabilities
2.1 Mathematics
Mathematics demands reasoning and a sophisticated understanding of complex information, making it one of the disciplines that place the highest premium on these cognitive abilities. There is significant academic interest in enhancing the mathematical capabilities of LLMs (Lu et al., 2022b). In the pursuit of creating an education system based on LLMs, our goal is for it to possess the ability to address various aspects of mathematical problems, whether they involve basic numerical computations, intricate reasoning, or problems integrating multi-modal information. In this section, we summarize the current developments in the mathematical capabilities of LLMs, focusing on four aspects: fundamental numerical computations, complex reasoning, the handling of multi-modal problem-solving, and mathematical proof.
2.1.1 Basic Arithmetic Problems
In human learning mathematics, foundational mathematical operations serve as the basis for addressing more advanced problems. Recently, significant scholarly attention has focused on augmenting LLMs’ proficiency in this domain (Patel et al., 2021; Wang et al., 2017; Zhang et al., 2023a; Geva et al., 2020). Most models try to fine-tune a pre-trained LLM on high-quality mathematics datasets. Given the robust comprehension of human language and notable textual reasoning abilities exhibited by LLMs, it is natural for many to assume that LLMs should effortlessly handle fundamental mathematical problems. However, the reality proves otherwise. Yuan et al. (2023) pointed out that ChatGPT and GPT-4 perform well in addition and subtraction operations, but their accuracy decreases when dealing with multiplication involving larger numerical values. This limitation arises because the language models do not access a calculator during the computations. Yang et al. (2023c) and Liu and Low (2023) proposed to fine-tune LLMs on high-quality datasets and found that even small models could avoid this multi-digit problem. However, it is not feasible to fine-tune each LLM individually. How to prevent such issues during the pre-training phase of language models remains an open question.
2.1.2 Challenging Mathematics Problems
Despite occasional errors in basic mathematical operations, the expectation for LLMs to solve more complex problems remains high, and the LLMs’ capabilities in this area continue to develop. The ability to handle college-level mathematics is particularly beneficial for senior students’ learning, offering significant assistance in understanding challenging concepts. In recent years, more and more work has been proposed to develop this ability (Luo et al., 2023a; Yu et al., 2023; Lewkowycz et al., 2022; Wang et al., 2023a). Wang et al. (2023a) introduced a benchmark SCIBENCH, which contains collegiate-level scientific problems from mathematics, chemistry, and physics textbooks. Luo et al. (2023a) tried to enhance the mathematical reasoning abilities of Llama-2 by applying Reinforcement Learning from Evol-Instruct Feedback (RLEIF). Zhou et al. (2023) tried to enhance GPT-4’s mathematical ability by encouraging it to use code to self-verify its answers. This approach leads to a significant improvement in the zero-shot accuracy of mathematical problem-solving.
2.1.3 Problems Involving Multi-Modal Information
Multi-modal inputs like geometric shapes in math problems require LLMs to understand text and image information for solutions. Research in multi-modal LLMs for mathematical reasoning is emerging (Lu et al., 2023; Chen et al., 2022b; Peng et al., 2023b). This type of task poses a high requirement for the training data’s formation and quality. Chen et al. (2022b) introduced a Unified Geometry problem benchmark combining calculation and proving tasks. Based on this dataset, the study presented a framework capable of simultaneously solving calculations and proving tasks through a sequence generation approach. Lu et al. (2023) proposed MATHVISTA, a benchmark for diverse mathematical and visual challenges. Frameworks for this task usually contain separate encoders for different modalities. For geometric problems, Zhang et al. (2023b) converted diagrams into text clauses, using a convolutional neural network and a language model for encoding and a GRU-based framework for answer generation. Additionally, Lu et al. (2022a) presented the Tabular Math Word Problems (TABWP) dataset, requiring textual and tabular data reasoning, and introduced PROMPTPG, a policy gradient-based selector for training and prompt construction for test samples.
2.1.4 Mathematical Proof
Unlike other types of mathematical problems where LLMs primarily focu on reasoning and providing answers, LLMs’ role in mathematical proofs emphasizes the integration with proof assistants such as Coq (Barras et al., 1997), Isabelle (Nipkow et al., 2002), and Lean (de Moura et al., 2015). These proof assistants correspond to specific programming languages, requiring users to input proofs in these languages, after which the assistant can verify the proof’s correctness.
Many LLM-based methods are proposed to help theorem proving (Jiang et al., 2023; Liu et al., 2023b; Xiong et al., 2023). Based on these proof assistants, there are two main approaches to utilizing LLMs for mathematical proofs. The first approach is formal proof search, exemplified by models like GPT-f (Polu and Sutskever, 2020), which involves prompting LLMs to produce the next proof step(also called ’tactic’ in proof assistants) based on the current proof state and some optional context. Following, Thor (Jiang et al., 2022b) was further proposed to help select the premise for theorem proving. Yang et al. (2023b) introduced an open-source framework named LeanDojo based on the Lean proof assistant. This framework comprises data, toolkits, models, and benchmarks, and it has led to the development of ReProver (Retrieval-Augmented Prover), which enhances proof accuracy by using retrieval methods to extract premises for LLMs to base their mathematical proofs.
The second is natural proof translation, also known as autoformalization, represented by models like DSP (Jiang et al., 2022a), which converts math proofs written in natural language into formalized versions. This approach primarily addresses the issue of insufficient data for formal mathematical proofs. By utilizing autoformalization, a significant increase in this type of data can be achieved. Consequently, this approach enhances the proof-generating capabilities of neural provers that have been fine-tuned on this expanded dataset. The strategy leverages the transformation of existing proofs into a formal structure, thereby enriching the training material for machine learning models in the field of mathematical proof generation. For example, Wu et al. (2022) employed LLMs for autoformalization, transforming mathematical proofs and problems expressed in natural language into formal specifications and proofs in Isabelle language. The generated data was used to train a neural theorem prover, enhancing the effectiveness of the original prover. Follow it, Cunningham et al. (2023) utilized an encoder-decoder framework based on the universal transformer architecture, converting both problem statements and mathematical proofs written in LaTeX into the language of the Coq interactive prover.
In mathematical education, proof problems are indispensable. Currently, LLMs for math proof primarily operate in the form of interactive theorem proving. In this approach, LLMs complete proofs by interacting with software proof assistants. To realize completely automated theorem proving with LLMs, it is essential that these models possess not only strong reasoning skills but also the capability to formalize concepts effectively. There is no room for hallucination in mathematical proofs, which poses a formidable challenge for LLMs.
2.2 Writing
Writing proficiency is crucial for LLMs, underpinning their ability to comprehend inputs deeply and produce semantically and syntactically accurate outputs (Dong et al., 2022; Chang et al., 2023). In education, the writing capability of LLMs holds the potential to transform how writing is taught. They can assist in content creation, simplify complex topics for students, and offer personalized educational materials. In this part, we dive into LLM’s writing capability on two education-related tasks: text summarization and grammatical error correction.
2.2.1 Text Summarization
Text summarization, which condenses long texts into summaries, challenges LLMs to understand diverse content like news and cross-lingual texts. It is evident that traditional fine-tuning methods are less effective with the advent of more advanced LLMs (Pu et al., 2023; Liu et al., 2022). Pu et al. (2023) and Liang et al. (2022) showed LLMs like ChatGPT initially lag behind fine-tuned models like T5 (Raffel et al., 2020) and BART (Lewis et al., 2019) in ROUGE scores (Lin, 2004) for text summarization. However, when human judges evaluate overall quality, LLMs outperform fine-tuned models and even standard human summaries, superior in aspects like factual consistency, fluency, and diversity. This discovery underscores the limitations of traditional evaluation methods and suggests a need for new paradigms to guide summarization tasks in the LLM era. For example, BRIO (Liu et al., 2022) implemented a ranking task to foster more diverse summarizations. Furthermore, Liu et al. (2023e) utilized a GPT model based on BRIO to directly generate training data to guide the learning process of other models, which is similar to the process of RLHF (Stiennon et al., 2020).
2.2.2 Grammatical Error Correction
We are well aware of the remarkable capability of LLMs to generate fluent and coherent conversations. However, from an educational perspective, the importance of producing grammatically correct dialogues cannot be overstated, especially for students learning a new language. The correctness of grammar in conversations plays a vital role in language acquisition, providing students with reliable examples to emulate and learn from. Numerous studies have evaluated the effectiveness of LLMs in grammatical error correction. Several works (Wu et al., 2023; Fang et al., 2023; Mohammed et al., 2023) first evaluated the error correction performance of closed-source LLMs such as ChatGPT. Although there exists a pronounced gap between ChatGPT and the previous SOTA (Omelianchuk et al., 2020; Grammarly, 2023) models on the overall F0.5 metric, closer analysis shows that ChatGPT underperforms other models in terms of precision but far exceeds other models in terms of recall. That said, LLMs like ChatGPT are good at error detection. A detailed manual analysis of ChatGPT’s outputs revealed that, in most cases, it maintained grammar accuracy better than the previous methods. However, it often overcorrects sentences to increase diversity and fluency, resulting in a decrease in precision. Nevertheless, it demonstrates that LLMs are adept at generating diverse and accurate content. Some studies have attempted to explore open-source LLMs, such as CoEdit (Raheja et al., 2023), which covered multiple text editing tasks (including GEC) by fine-tuning LLMs to integrate the capabilities brought by these tasks. GrammarGPT (Fan et al., 2023), on the other hand, asked ChatGPT to generate grammatically incorrect sentences to guide the training of other models. Collectively, these studies underscore the prowess of LLMs in generating correct text.
Leveraging LLMs’ proficiency in text summarization and grammatical error correction can significantly benefit education. Their capability to condense complex material into concise summaries facilitates efficient learning, while error correction tools help improve students’ writing and language skills. However, critical challenges need to be resolved to integrate these writing-related capabilities to help education. It becomes evident that more refined evaluation metrics and task-specific optimizations are essential for LLMs.
2.3 Programming
Programming is a process of writing code and correcting code if unexpected results are obtained. Incorporating LLMs in programming education is reshaping the future of AI-assisted programming learning. LLMs could play multiple roles: as instructors providing guidance, as teaching assistants offering personalized tutorials, and as collaborative coding partners. Studies like (Ma et al., 2023a) demonstrated improved performance (17%) and efficiency (13%) among programming novices using LLM-based assistants. Research from (Phung et al., 2023) focused on programming education tasks and benchmarks like (Fu et al., 2023) and (Ding et al., 2023), which were used to evaluate the effectiveness of LLMs. This section mainly discusses the LLMs’ coding development from two perspectives: code writing and code refinement, corresponding to the two stages in programming.
2.3.1 Code Writing
Unlike natural language tasks, generating code requires a more rigorous token syntax and places higher demands on the training stage. A common method to improve LLMs’ performance in generating code is to train or fine-tune them on extensive code datasets (Chen et al., 2021a; Nijkamp et al., 2023). WizardCoder (Luo et al., 2023b) introduced the Evol-Instruct (Xu et al., 2023) method to generate more complex and diverse instruction datasets of code-related tasks. To emulate the iterative process of humans repeatedly modifying and reviewing code, InCoder (Fried et al., 2023) utilized bidirectional encoding instead of left-to-right encoding. In addition to next-token prediction, training or fine-tuning code-aimed LLMs on additional code-related tasks could enhance their programming capabilities. LLMs first learn language patterns and representations from a large amount of text data through unsupervised learning. Then, they could be fine-tuned on labeled code tasks, allowing them to learn more targeted code representations and gain a deeper understanding of code structure and semantics based on the provided labels. CodeT5+ (Wang et al., 2023b) introduced the concepts of unimodal and bimodal alignment, increasing the model’s adaptability to function in different modes for various downstream tasks. During the bimodal alignment phase, the model synchronizes the representations of text-code pairs using multiple tasks, which improves its ability to understand and generate content across different modalities. CodeLlama (Rozière et al., 2023) also applied the multi-task objectives, including autoregression and causal infilling prediction, which achieves better performance among open models. MFTCoder (Liu et al., 2023a) utilized the Multi-Task Learning (MTL) technique and incorporated a training loss computation algorithm to alleviate the instability and imbalance of multi-task training.
2.3.2 Code Refinement
In most cases, LLMs could not generate the correct code at once. We could enable LLMs to generate a code sketch (either actual code or pseudocode) and utilize various methods to guide the model to modify and refine the code. By leveraging the inherent code correction ability of LLMs, the overall precision and quality of the code could be significantly enhanced. Liventsev et al. (2023) constructed a pipeline: Synthesize, Execute, Instruct, Debug, and Rank (SEIDR). It first generates multiple different codes and undergoes the process of code filtering and debugging, ultimately selecting the best code among them. According to Magister et al. (2022), teaching an LLM to debug its program draft via few-shot demonstrations could improve the performance on code generation tasks. Another method for LLM debug is generating unit tests by LLM itself and checking its code (Chen et al., 2022a). By mimicking the human coding process, LLM’s programming ability is greatly enhanced. However, these methods lead to an increased number of calls to LLMs, resulting in a significant increase in inference time.
In coding education, the guidance given by current LLMs still falls short of the level offered by human instructors. Firstly, there is significant room for improvement in the coding abilities of LLMs. When it comes to generating complex algorithms, LLMs’ performance rapidly declines compared to the functionality achieved by humans (Chen et al., 2021b). Additionally, due to the lack of real-world data, LLMs struggle to learn the intermediate thinking process of code writing, making it difficult for them to provide relevant explanations and instructions to beginners. As a result, the use of LLMs in programming education still needs to be improved, especially in terms of interpretability.
2.4 Reasoning
The reasoning capability of LLMs offers significant potential for educational use. LLMs can act as advanced tools, enhancing students’ cognitive processes, providing personalized mentorship, and offering tailored learning support. Consequently, these models can bolster students’ problem-solving skills and promote holistic cognitive development in an educational setting. This section reviews LLMs’ general reasoning ability development strategies.
2.4.1 Supervised Fine-tuning for Reasoning
Prior researches on the LLMs’ reasoning ability are primarily based on the fully supervised fine-tuning method. For instance, Rajani et al. (2019) effectively fine-tuned a pre-trained GPT model to generate rationales for predictions based on the CoS-E dataset (Talmor et al., 2018), enhancing performance in commonsense question-answering tasks. However, the success of fine-tuning mainly depends on the availability of a dataset with explicit reasoning, which could be challenging to obtain. Furthermore, the inferences drawn from fine-tuning are mainly limited to specific domain datasets, relying more on the content within the fine-tuning data than on the LLM’s inherent reasoning abilities.
2.4.2 Prompt Engineering for Reasoning
Efforts have been made in recent research to tackle the constraints inherent in the fine-tuning process of LLMs. These fine-tuning methods tend to overfit specific dataset distributions, reducing their effectiveness on more diverse datasets. To mitigate this issue, numerous studies have developed strategies that utilize the robust reasoning abilities of LLMs, taking advantage of their extensive pre-trained parameters. One such strategy involves guiding LLMs to generate inference and reasoning processes through demonstrations or prompts. For example, Wei et al. (2022) introduced the ”Chain of Thought” (CoT) method, which employed natural language reasoning steps as prompts to enhance the reasoning performance of LLMs. Similarly, Wang et al. (2022) proposed a self-consistency strategy involving sampling a variety of reasoning paths and selecting the most consistent answer, which has been shown to improve model performance. Addressing the constraint of fixed manual annotation demonstrations, Diao et al. (2023) introduced an active selection approach. This method dynamically identified the most effective demonstrations relevant to task requirements from a pool of queries. Furthermore, Zhou et al. (2022) proposed a prompting strategy that decomposed complex problems into simpler sub-problems. This approach encouraged sequential problem-solving and could potentially enhance the performance of LLMs further.
2.4.3 Hybrid Methodologies for Reasoning
Despite the success of prompt engineering in improving the performance of LLMs, it does not boost their core reasoning skills by not changing the model parameters. To address this limitation, new approaches combining fine-tuning with prompt engineering have been developed, aiming to enhance the model’s reasoning abilities and responsiveness to prompts. One approach employs LLMs in a pedagogical capacity to instruct smaller student models. Studies by Ho et al. (2022) and Magister et al. (2022) demonstrated significant performance gains across various datasets through fine-tuning student models based on the reasoning outputs of larger LLMs, even for smaller-scale models. Moreover, Zelikman et al. (2022) reported substantial performance improvements across multiple datasets by generating step-by-step rationales and fine-tuning models based on correct answers, thus facilitating model learning from its reasoning.
In conclusion, the discussed methods have significantly advanced the reasoning capability of LLMs, optimally utilizing their unique features for diverse reasoning tasks. This enhancement can greatly benefit educational applications. However, it’s crucial to recognize the limitations. As underscored by Valmeekam et al. (2022) and Ruis et al. (2022), LLMs struggle with complex reasoning tasks and those requiring implicit expressions. For example, LLMs can struggle with complex reasoning scenarios, leading to a notable decrease in performance. This is particularly relevant in educational contexts, where incorrect problem-solving modeled by LLMs could misguide students and lead to misunderstandings or flawed comprehension. Thus, despite LLMs’ immense potential in education, their limitations must be carefully considered to ensure they facilitate rather than obstruct learning.
Table 1: Overview of LLMs’ performance on foundational education-related capabilities.
Models Reference Mathematics Writing Programming Reasoning Knowledge-based QA General
(GSM8K) (OpenCompass) (HumanEval) (HelloSwag) (TruthfulQA) (C-Eval)
(Pass@1) (Avg Score) (Pass@1) (Acc) (MC2) (Avg Score)
GPT-4 (OpenAI, 2023b) 92.00 62.00 67.00 91.40 59.00 68.70
ChatGPT (OpenAI, 2023a) 57.10 48.60 48.10 79.50 47.00 54.40
TigerBot-70B-Chat-V2(70B) (TigerResearch, 2023) 54.36 61.30 30.50 82.83 75.40 -
LLaMA2(70B) (Touvron et al., 2023b) 60.27 51.60 29.90 82.30 56.18 55.20
LLaMA(65B) (Touvron et al., 2023a) 43.37 47.10 23.70 82.30 55.09 38.80
Yi(34B) (lingyiwanwu, 2023) 50.64 48.90 26.20 82.00 56.23 81.40
Vicuna(33B) (TheVicunaTeam, 2023) 13.72 44.90 15.20 83.00 56.16 39.80
WizardLM(30B) (Xu et al., 2023) 34.42 - 26.08 76.32 49.14 -
Moss(16B) (FudanUniversity, 2023) 6.90 39.00 - 55.80 49.00 33.10
Qwen(14B) (Bai et al., 2023) 58.98 52.7 43.90 80.20 49.43 72.10
Baichuan2(13B) (Yang et al., 2023a) 55.30 51.50 17.07 66.90 48.98 40.00
Alpaca(7B) (Taori et al., 2023) 0.15 39.50 9.10 75.71 36.28 29.90
ChatGLM3(6B) (Zeng et al., 2022) 72.30 43.10 44.50 76.50 - 69.00
2.5 Knowledge-based Question Answering
In the context of Knowledge-based Question Answering using LLM, the user presents a question to LLM, and LLM leverages knowledge-based methods and responds with the corresponding answer. Previous work by Ren et al. (2023) showed that LLMs have an inaccurate perception of factual boundaries and often exhibit overconfidence. Many studies have explored and utilized external knowledge from open-world and domain-specific databases to enhance the knowledge base of these LLMs.
2.5.1 Open-domain QA
Open-domain question answering requires LLMs to accurately determine the reliability of information in the open world and craft their responses based on that understanding. Jiang et al. (2021) evaluated the accuracy of LLM responses to a particular question from the perspective of calibration. Through experiments, the researchers discovered that models such as T5, BERT, and GPT-2 are not well-calibrated in QA tasks. While suggesting that incorporating calibration-related methods into the fine-tuning process can effectively enhance performance in QA tasks, it is evident that solely pre-trained language models still face significant challenges in open-domain tasks. To overcome this challenge, many works tried to add additional information to help the LLMs answer correctly (Khandelwal et al., 2019; Guu et al., 2020; Borgeaud et al., 2022). Lazaridou et al. (2022) employed information gathered from web searches as prompt input for LLM, conditioning it to generate answers to questions. This approach effectively enables LLM to use open-world information to answer questions.
The development of LLM-based open-domain question answering highlights significant challenges, particularly in dealing with hallucinations. In the context of establishing an LLM-based education system, this issue becomes more critical for providing seemingly correct yet incorrect answers, leading to misleading students. Drawing insights from approaches that introduce additional information from sources such as the web or textbooks can offer valuable lessons for the development of an LLM-based education system.
2.5.2 Domain-specific QA
Although LLMs are trained on vast corpora, they may still exhibit gaps in understanding specific domains. Typically, there are dedicated knowledge repositories for professional content that consolidate domain-specific knowledge, such as MedlinePlus [^2] https://medlineplus.gov, GeeksforGeeks [^3] https://www.geeksforgeeks.org, etc. Choi et al. (2023) utilized an external knowledge base to generate a set of question-answer pairs and then employed fine-tuning to transfer financial knowledge to LLMs, significantly improving financial question-answering tasks. Another common approach is to leverage the in-context learning capability of LLMs by incorporating retrieved knowledge from the knowledge base into prompts. Peng et al. (2023a) used text embeddings for retrieval, followed by LLM’s automatic feature extraction in order to achieve higher accuracy in pest identification tasks. zhang et al. (2023d) utilized K-nearest neighbors (KNN) to search for the most similar K records from an accounting database, serving as k-shot examples and greatly improving accounting efficiency. There are also works that train and improve the retriever encoder (Zhang et al., 2023c), as well as distill and refine the data in the database (Jeronymo et al., 2023). Such retrieval frameworks have lower costs and can be more flexible in applications across different domains. Liu et al. (2023c) built upon a retrieval-enhanced framework and provided several plug-and-play modules, enabling its application in various professional fields.
By integrating Information Retrieval (IR) systems, LLMs can enhance their capabilities with professional knowledge, gaining valuable and precise supplemental information. Furthermore, according to Ren et al. (2023), retrieval augmentation can also be employed to improve LLMs’ ability to perceive facts within the boundaries of their legal knowledge, mitigating the issue of hallucinations. During the education process, different majors or courses involve different professional content. Applying external knowledge repositories as an enhancement mechanism can provide more accurate guidance in domain-specific contexts and mitigate the issues caused by misleading information. Therefore, domain-specific question-answering ability is crucial for developing an LLM-based education system.
2.6 Overall Development Status
We select representative benchmarks to assess the current development of LLMs across education-related capabilities. Specifically, we mainly collect the results from three sources: Huggingface [^4] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass [^5] https://opencompass.org.cn/leaderboard-llm and C-Eval [^6] https://cevalbenchmark.com/static/leaderboard.html. The formal two are comprehensive leaderboards. C-Eval is a Chinese evaluation suite for foundation models spanning 52 diverse disciplines. We collect performance data from popular general LLMs on these benchmarks, and the compiled results are presented in Table 1, where one can observe that:
  • It is hard for a single LLM to be superior across all capabilities. GPT-4 exhibits the best overall performance among current LLMs. However, LLMs still lag significantly behind humans in some crucial abilities, such as achieving 94% accuracy in TruthfulQA (Lin et al., 2021), a performance level far superior to the optimal results of LLMs.
  • Most LLMs display considerable variation in developing these skills. While certain models (such as Alpaca and Yi) might excel in text comprehension tasks, their effectiveness often diminishes in areas requiring deep understanding and reasoning, like Mathematics and Programming. This reveals the substantial challenges in building a unified education-focused LLM since it may fail in certain areas.
((a))
((b))
Figure 3: Two frameworks towards LLM-based educational framework. (a) is the unified framework where one LLM model handles all the education-related questions. (b) is the MoE framework where the LLM controller is responsible for task distribution, and the corresponding expert model finishes the given task.
3 Potential of LLM-based Education System
LLMs can potentially transform online education by understanding a wide range of student questions, similar to human teachers. They aim to provide support across different subjects and skill levels. With the latest developments in LLMs, we suggest two approaches for creating LLM-based education systems. The first involves training a comprehensive LLM that can handle questions from various subjects. The second approach uses a mixture-of-experts (MoE) framework, integrating specialized models to support the system with an LLM controller to manage interactive dialogs with students.
3.1 Unified Approach
The most straightforward idea for establishing an LLM-based education system is to train a language model capable of answering students’ questions across all subjects. As shown in Figure 3 (a), the foundational capabilities are included in the unified LLM, and the student can directly communicate with it and ask questions.
Research on whether LLMs can handle educational tasks has been underway. Wang and Demszky (2023) introduced three teacher coaching tasks for generative AI: (A) scoring transcript segments using classroom observation instruments, (B) identifying highlights and missed opportunities for effective instructional strategies, and (C) offering actionable suggestions to encourage more student reasoning. And evaluated by human teachers, ChatGPT on these tasks for elementary math classroom transcripts generates responses that are relevant to improving instruction, but they are often not novel or insightful. Beyond that, Phung et al. (2023) assessed the programming education ability of ChatGPT and GPT-4 by comparing them with human tutors. The result shows that GPT-4 performs way better than ChatGPT, even close to human tutors in some scenarios, while it also highlights some situations in which GPT-4 struggles. In particular, for the grading feedback and task creation scenarios that have a substantial gap in the performance of GPT-4 compared to that of human tutors.
3.2 MoE Approach
Section 2 reviewed the current development of LLMs across various capabilities. Unfortunately, despite the existence of comprehensive language models, such as GPT-4, these models often exhibit notable deficiencies in certain abilities. This situation poses a challenge, indicating that relying solely on an LLM itself for educational guidance involving all these capabilities is currently a difficult task. Yet, LLMs can achieve excellent results through fine-tuning individual capabilities, and their ability to comprehend human language is exceptionally strong. Therefore, we can aggregate models with distinct capabilities using a mixture-of-experts approach. By establishing an LLM-based controller for language interaction and task assignment with students, a currently feasible education system can be generated.
An education framework implemented with a mixture-of-experts (MoE) approach is illustrated in Figure 3(b), consisting of multiple models that excel in individual capabilities (not necessarily LLMs) and an LLM controller. The controller is mainly responsible for three tasks: (1) Understand the student’s request and decide which specific area or areas the request is about. (2) Re-form the request to fit the input of the specific areas’ expert models. (3) Aggregate the output of the related experts and generate the final response to the student.
The advantage of this approach is that training is less challenging, requiring training on individual capability-related datasets to obtain multiple models, each superior in a specific capability. However, the drawback lies in potential understanding errors during inference due to interactions. This includes errors introduced by the LLM controller when assigning tasks based on student input and integrating outputs from various experts. Currently, this approach might be a realistic way of achieving the LLM-based educational assistant system.
4 Future Directions
Recently, more and more researchers have been trying to apply LLMs to handle education tasks, such as course design, student evaluation, lesson plan design, and others. Nevertheless, there are still many challenges and opportunities.
4.1 Interdisciplinary Reasoning Ability
Students may encounter interdisciplinary reasoning problems during real-world learning, requiring the education system to integrate multiple capabilities to formulate responses. As illustrated in Figure 1, the student intends to write a program to solve a mathematical problem, and the model needs first to comprehend the mathematical problem, devise a solution, and then generate the code. This process necessitates the model to synthesize both mathematical and programming capabilities.
However, there is currently limited research in the integration of multiple interdisciplinary capabilities for LLMs at this stage, including both datasets and algorithms. Boyko et al. (2023) examined how LLMs augment scientific inquiry, code development, scientific writing process, etc., and they propose that LLMs can foster interdisciplinary work by bridging knowledge gaps across scientific fields. However, they mainly discuss the LLMs’ ability to help researchers’ interdisciplinary collaboration instead of their ability to answer interdisciplinary questions. Cultivating an LLM to obtain this ability would help to develop a unified education system, which is an essential research direction.
4.2 Student Modeling
Establishing an LLM-based education system allows students to articulate their personalized needs through dialogs. Through such conversations, we can extract or infer personalized features about students, such as their current mastery of topics and preferences in learning styles.
Besides modeling students from conversations, some researches (Aher et al., 2023; Argyle et al., 2023) have shown that LLMs have certain abilities in simulating humans and generating human samples. Applying to education, this ability indicates a potential for LLM-based student simulation. In this way, for the students with few interaction records, the LLM-based simulator could generate more samples and provide data to help the expert model better understand the student. It could help human teachers develop teaching skills better.
4.3 Multi-modal Education
In education, multi-modal information is common, like geometry problems combining text and images or textbook concepts with illustrations. Building a general intelligent education system requires handling such multi-modal data. Thus, developing a multi-modal educational model is a promising future direction.
Notably, the development of multi-modal LLMs is rapidly advancing (Ye et al., 2023; Du et al., 2022). Different kinds of architectures and pre-train tasks are proposed (Du et al., 2022). However, the education domain often exhibits unique distribution characteristics in multi-modal information. Therefore, it may place higher demands on the model’s multi-modal reasoning capability.
5 Conclusion
In this paper, we presented an overview of the development of the LLM-based education system. We first reviewed the important development of LLMs’ education-related abilities. Then, we analyzed the potential of it and proposed two different ways of building such a system. We also highlighted the future directions that are worth working on. We hope this survey provides some insight into future research in this direction.
Acknowledgments
The SJTU team is partially supported by National Natural Science Foundation of China (62177033).